OpenStudy (anonymous):

Optimization question: Find the dimensions of a rectangle of largest area that can be inscribed in the circle x^2 + y^2=9

OpenStudy (anonymous):

what is the area?

OpenStudy (amistre64):

its a square.... just to let you know ;)

OpenStudy (anonymous):

can you please help me out and tell me how to get the solution?

OpenStudy (amistre64):

the solution to this is to just use 1/4 of the circle the part that lies in the first qudrant

OpenStudy (amistre64):

we know the are of the 'rectangle' = xy right? and our 'line' is full of x and y options; so lets use the equation to define x or y in terms of the other...

OpenStudy (anonymous):

yes i got y= square root of 9-x^2

OpenStudy (amistre64):

x^2 + y^2 = 9 x^2 = 9 - y^2 x = sqrt(9-y^2) ; use this value of 'x' in the rectangle area equation

OpenStudy (amistre64):

y = ... is fine to; doesnt matter lol

OpenStudy (anonymous):

lol okay then just substitute inside the equation A=xy

OpenStudy (amistre64):

yes; then derive :)

OpenStudy (anonymous):

okay thank you so much

OpenStudy (amistre64):

just to let you know; the optimum is at 45 degrees; or rather when y = 9sin(45) :)

OpenStudy (anonymous):

okay thank you